OpenStudy (cliffordtapp):

victor needs to calculate the number of miles per gallon his car averages. his cars gas tank hold 15 gallons of gas. determine the average number of miles his car gets per gallon of gas. lets let x=number of miles traveled a=average miles per gallon

OpenStudy (jameshorton):

cliff wat school u go to

OpenStudy (cliffordtapp):

iowa high

OpenStudy (jameshorton):

wat grade

OpenStudy (jameshorton):

X

OpenStudy (cliffordtapp):

tenth

OpenStudy (cliffordtapp):

why you trying to avoid the question

OpenStudy (jameshorton):

the answer is X

OpenStudy (cliffordtapp):

that don't make NO SENSE you need to write an equation

OpenStudy (jameshorton):

you doing algebra 1 or just algebra

OpenStudy (cliffordtapp):

algebra 2

OpenStudy (anonymous):

i can help you

OpenStudy (cliffordtapp):

would the answer be x=a/15

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

a = x/15 if he drives until his tank becomes empty

OpenStudy (anonymous):

do u want me to go through the steps though

OpenStudy (cliffordtapp):

no you don't have to go through the steps

OpenStudy (anonymous):

ok tag me if u need furthur help :)

OpenStudy (cliffordtapp):

so the answer would be a=x/15

OpenStudy (welshfella):

the answer is a = x/15 not x = a/15

OpenStudy (welshfella):

yea

OpenStudy (cliffordtapp):

thanks it was right

OpenStudy (welshfella):

yw